OpenStudy (anonymous):

cosec x= 1/2 (x)+1 has root 0

OpenStudy (anonymous):

Remember -1 <= sin x <= 1 Therefore -1 <= cosec x <=1 Now try :)

OpenStudy (anonymous):

Still not sure..

OpenStudy (anonymous):

One more question. This is your question , right \[cosec x= {1 \over 2 }(x)+1\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I think i solved it

OpenStudy (anonymous):

OK. Just put instead of cosec x , 1/2(x) +1

OpenStudy (anonymous):

Ok @NotSObright , let's see what you have done :P

OpenStudy (anonymous):

Say y1=1/2x+1 y2=csc(x) y1(0)=1 y1(Pi/2)=Pi/4+1 y2(0)=infinty y2(Pi/2)=1 Since it can be proven using differentiation y1 is inceasing and y2 is decreasing on 0,pi/2 hence in order to attain boundary values graphs must cross each other

OpenStudy (anonymous):

So what about the 0.5 and 1?

OpenStudy (anonymous):

sorry i didn't see that u must show at 0.5 y1<y2 and at 1 y1>y2 u might wanna use a calculator

OpenStudy (anonymous):

But how?

OpenStudy (anonymous):